OpenStudy (anonymous):

What is the axis of symmetry for f(x)=2x^2-8x+8? x=-2 x=-3 x=3 x=2

OpenStudy (anonymous):

axis of symmetry : x = -b/(2a)

OpenStudy (helder_edwin):

try completing the square. do know how?

OpenStudy (helder_edwin):

once u have the equation in the form \[ y-k=A(x-h)^2 \] u can get all the information u might want

OpenStudy (anonymous):

h will always = -b/2a

OpenStudy (anonymous):

so it would be 8/2(2) which would be reduced to 2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright thank you.